40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 23 17:30:33 2012, jamie, nic, jenne, raji, manasa, Update, Alignment, Initial attempts to fix IFO alignment 
    Reply  Tue Oct 23 17:41:20 2012, Manasa, Update, Alignment, Power supply at OMC removed 
       Reply  Tue Oct 23 18:21:21 2012, Jenne, Update, Alignment, Power supply at OMC removed 
    Reply  Tue Oct 23 18:12:18 2012, Jenne, Update, Alignment, Tip tilt wires - the truth 
Message ID: 7601     Entry time: Tue Oct 23 18:12:18 2012     In reply to: 7599
Author: Jenne 
Type: Update 
Category: Alignment 
Subject: Tip tilt wires - the truth 

Quote:

At this point I cried foul.  This is not an acceptable situation.  Very little stimulation to the tip-tilts can repoint the beam inside the PR cavity.

Steve says that the TT weights, which will attach to the base of the TT mirror mounts and should help keep the mirrors vertical and not hysteretic, are being baked now and should be available tomorrow. 

 We also wrote down the serial numbers (top center of each TT, inscribed by hand) for what tip tilt is installed where.  I then went through the elog to determine which TT was suspended with what kind of wire (thick or thin).  Summary: all installed tip tilts have thick wire, 0.0036" diameter.

As noted in elog 3295, we had found that there was similar hysteresis whether we used the thick or the thin wire, so we had decided not to go back and re-suspend every optic.

Also, since we will redo the pitch balance tomorrow with the new hardware tomorrow, I think we should put in the new LaserOptik mirrors at the same time.  We have not yet gotten phase maps of them, but we might as well do this rebalancing once, rather than twice.

 

As-installed tip tilt list
Serial number Installed as Wire thickness Notes, elog reference
001 SR 3 0.0036" See elog 3437
002 SR 2 0.0036" See elog 3295
003 PR 2 0.0036" No elog, but inferred since there were 4 with thick wire, and #004 is the thin wire one.  Elog 3437 has notes on the 4 thick, 1 thin situation.
004 spare, dirty originally 0.0017", but looks redone with thicker wire See elog 3295
005 PR 3 0.0036" Was supposed to be spare according to elog 3437, but was installed.  See elog 3437

 

ELOG V3.1.3-